OpenStudy (abbyrose0105):

(1 + 2i) • (5 − 3i) How I solve this? I thought that I would first use FOIL to distribute and then add the like terms but that was wrong...

OpenStudy (anonymous):

What do you mean it was wrong? Can you show me your steps?

OpenStudy (abbyrose0105):

Simplify: (1+2i) * (5-3i) 5-3i+10i-6i =5+1

OpenStudy (anonymous):

http://www.freemathhelp.com/using-foil.html If you need help using the foil method heres a great reference

OpenStudy (abbyrose0105):

@bradely I have 4 options and that isn't one of them

OpenStudy (anonymous):

In the step were you multiply the last term in parenthesis when you multiply two terms with a variable you square it, so 6i should have been 6i^2. Heres your final trinomial. 6i^2 + 7i +5

OpenStudy (anonymous):

-6i^2 +7i+5 sorry forgot the negative :)

OpenStudy (abbyrose0105):

These are my four options: 2 − 15i 3 + 2i 5 − 6i 11 + 7i

OpenStudy (anonymous):

11 +7i is correct!

OpenStudy (abbyrose0105):

Thanks guys!

OpenStudy (bradely):

(1 + 2i) • (5 − 3i) =1•(5 − 3i)+2i•(5 − 3i) =5-3i+10i-(2i)(3i) =5+7i-6i^2 where i^2 =-1 =5+7i-6(-1) =5+7i+6 =11+7i Option 4 is correct Further help http://www.mathskey.com/question2answer/ask